# Boosting quantum annealing performance through direct polynomial unconstrained binary optimization

### v2

@@ IV.1.3 PUBO-to-QUBO reduction
To assess the annealing performance gains of PUBO over QUBO formulations when solving 3-SAT instances, we compare the minimum energy gap of the PUBO formulation with that of the equivalent QUBO reduction. As mentioned before, finding an optimal QUBO reduction for a given PUBO problem is in general $\mathsf{NP}$-hard. For the small system sizes considered here, we employ the following brute-force algorithm: Replace the pair of variables that appears most frequently in the third-order terms of the PUBO cost function by a slack variable (see section III.1), then continue with the second-most-frequent pair of variables and so on, until all third-order terms have been replaced by quadratic ones with associated slack variables and constraints [51]. For each newly introduced ancilla variable $y=x_{i}x_{j}$, a penalty term $\propto(3y+x_{i}x_{j}-2x_{i}y-2x_{j}y)$ needs to be added to the cost Hamiltonian (see section III.1) and scaled by a suitable factor, thus ensuring that violating the penalty term in favor of the three-body interactions is never energetically favorable. This factor depends on the problem size and structure: In our numerics, we make the simple choice that for every three-body interaction term $a_{ijk}x_{i}x_{j}x_{k}$ that is to be reduced by the ancilla variable $x_{\mathrm{ancilla}}=x_{i}x_{j}$, we add $|a_{ijk}|+1$ to the strength of the corresponding penalty term. In section IV.2.5, we discuss how varying the strength of the penalty terms relative to the original cost Hamiltonian affects the overall annealing performance.
@@ IV.2.4 Dependence of the minimum energy gap on the QUBO penalty strength
To summarize, our example demonstrates that there is in general potential for optimization of the minimum energy gap in the QUBO reduction of a native PUBO problem by tuning the strength of the penalty terms. However, finding the optimal penalty strength generally requires solving the problem itself, which is prohibitively expensive for large exponentially hard problems. For our scaling analysis of the minimum gap, we have thus chosen the penalty strength in a problem-agnostic way as described in section IV.2.3. This heuristic choice is simple, incurs only moderate performance overhead, and reliably produces feasible solutions.
@@ V Implementing PUBO via multi-qubit interactions in quantum hardware
Figure 6: (a) The three-body interaction of the form $\mathinner{\hat{s}_{1}^{z}\hat{s}_{2}^{z}\hat{s}_{3}^{z}}$ can be decomposed into (a) four CNOT gates plus a single-qubit rotation. Each CNOT gate can be further decomposed into (b) one $R_{zz}$ gate of fixed angle and multiple single-qubit rotations. The quantum gates in the decompositions are defined as $R_{\alpha}(\theta)=\exp(-i\theta\hat{s}^{\alpha})$, with $\alpha\in\{x,y,z\}$, $R_{zz}(\theta)=\exp(-2i\theta\hat{s}^{z}\hat{s}^{z})$, and $H$ denoting the Hadamard gate.
@@ V.1 Native multi-qubit interactions and synthesis of three-body gates in quantum circuits
Several proposals exist to directly engineer three-body spin interactions of the type $\hat{s}_{i}^{z}\hat{s}_{j}^{z}\hat{s}_{k}^{z}$ ($ZZZ$ gate). For example, on trapped-ion hardware these include concurrent off-resonant driving of first and second sideband [54] (also relevant for other three-body interactions [55, 56, 57]) or a pulse sequence using squeezings and displacements of a phonon bus [58]. Another approach is via Floquet engineering of two-body interactions to obtain the three-body terms in a high-frequency expansion [59]. In addition, proposals for effective multi-body interactions on superconducting qubit architectures have been put forward [60].
Most of the available quantum computing architectures, however, implement natively only single- and two-qubit gates. These can realize higher-order interactions, required for PUBO formulations, with only polynomial overhead. An example is given in Fig. 6: A three-body term of the form $\hat{s}_{1}^{z}\hat{s}_{2}^{z}\hat{s}_{3}^{z}$ can be synthesized from four CNOT gates, plus a single-qubit rotation around the $z$-axis to obtain a variable strength [fig. 6(a)]. Since
@@
×$\times R_{y_{1}}\left(-\frac{\pi}{2}\right)R_{x_{1}}\left(-\frac{\pi}{2}\right)R_{x_{2}}\left(-\frac{\pi}{2}\right)H_{x_{1}}H_{x_{2}}$
×$\times R_{zz}\left(\frac{\pi}{2}\right)H_{x_{1}}H_{x_{2}}R_{y_{1}}\left(\frac{\pi}{2}\right)\,,$ (21)
where we used a two-qubit controlled rotation gate $R_{zz}$, Hadamard gates $H$ and single-qubit rotations around x and y-axis. We show the decomposition in fig. 6(b). The overall circuit for the three-body gate then uses four two-body gates of type $R_{zz}$ with fixed angle $\pi/2$ and 29 single-qubit rotations. In this decomposition, two adjacent single-qubit gates could be further combined into only one single-qubit rotation. Similar decompositions are possible with other two-qubit gates like the Mølmer–Sørensen (MS) type gate $R_{xx}$.
@@ V.2 PUBO speedup
As suggested by our numerical results in section IV, the minimal gap decreases more slowly with increasing system size for the PUBO formulation of certain 3-SAT problems than for the corresponding QUBO reduction, indicating that there exists a system size above which the PUBO formulation will always be favorable. Specifically. this is the case for 3-SAT instances generated by $\mathsf{toughSAT}$ and $\mathsf{uniquePT1}$, see fig. 3(b).
@@
Even in scenarios where PUBO and QUBO scale approximately the same, as is the case for the 3-SAT problems generated by $\mathsf{uniquePT1}$ shown in fig. 3(a), a constant PUBO speedup is still possible. We can use the estimated overhead of the three-body interactions from section V.1 together with the minimal gaps extracted numerically in section IV in order to estimate the potential gain in computing time when using the direct PUBO formulation according to eq. 8. In the example of fig. 6, executing the three-qubit gate involves four two-qubit gates and therefore takes about four times as long as a single two-qubit gate (neglecting single-qubit gate times). In the context of quantum annealing, this corresponds to a ratio of three-body to two-body interaction strength of $J^{(3)}/J^{(2)}\approx 1/4$. Using eq. 8 and the average values for the $\mathsf{uniquePT1}$ generator listed in table 1, we obtain $T_{Q}/T_{P}\approx$12.8$\mathinner{(\tilde{V}_{Q}/\tilde{V}_{P})}\mathrm{e}^{$0.024$\,N}$. In our numerics for small systems, we find a typical ratio of $\tilde{V}_{Q}/\tilde{V}_{P}\approx 0.5$, which only weakly depends on the system size. Our estimate therefore hints at a constant PUBO speedup in this regime; more precisely, an adiabatic sweep can be about six times faster if the PUBO instead of the QUBO formulation for this class of 3-SAT problems is used. If there is on top of that an (even small) exponential enhancement, as suggested by our numerical results, it will then dominate over a possible polynomial scaling of $\tilde{V}_{Q}/\tilde{V}_{P}$ at larger system sizes and further improve the performance of PUBO over QUBO. Note that the precise value of the estimated speedup depends on the choice of the strength of the driving Hamiltonian as well as the strength of the constraint terms in the QUBO reduction. Using a better heuristic for the latter could possibly reduce the calculated PUBO improvement. From our analyses in sections IV.2.4 and IV.2.5, where we found the used parameters to be close to optimal, we expect such a change to be only minor.
@@
For instances generated by $\mathsf{toughSAT}$, we see numerical evidence for a much stronger exponential scaling advantage of the PUBO formulation compared to the corresponding QUBO reduction: under the same assumptions as above, we find $T_{Q}/T_{P}\approx$0.81$\mathinner{(\tilde{V}_{Q}/\tilde{V}_{P})}\mathrm{e}^{$0.49$\,N}$. Already for small systems with $N=11$ classical variables, the adiabaticity time reduces by about two orders of magnitude in the PUBO formulation. That is, solving this family of exponentially hard 3-SAT problems via quantum annealing is expected to be two orders of magnitude faster when using the PUBO formulation over the conventional approach applying a QUBO reduction, and this estimate increases by another order of magnitude for every $\approx 4.7$ additional classical variables in the problem.
@@ VI Conclusion
In this work, we have presented a variety of combinatorial problems that are naturally phrased in terms of higher-order polynomial unconstrained binary optimization. Using the PUBO formulation directly instead of the corresponding QUBO reduction—currently the standard choice in quantum annealing—can have a number of advantages: (1) PUBO formulations avoid additional ancillary qubits, thus reducing the spatial resources required. (2) The ancillary qubits required by QUBO reductions are typically subject to consistency constraints, which require large energy scales—a central limiting factor in existing quantum annealing hardware. (3) As we have illustrated numerically at the example of the paradigmatic 3-SAT problem, PUBO formulations can have significantly larger minimum energy gaps, allowing for faster annealing sweeps. This gain outweighs the cost incurred by synthesizing the higher-order interactions through universal sets of single- and two-qubit gates. In particular, we find an exponential advantage of PUBO over QUBO in the scaling of the minimum energy gap with the system size for certain classes of problems (e.g., already for small $\mathsf{toughSAT}$ instances with $N=11$ variables, this suggests a potential speedup of about two orders of magnitude). Altogether, implementing optimization problems in quantum annealers as native PUBO rather than QUBO can thus lead to savings in terms of both spatial and temporal resources.
@@
Although in this work we have focused on the paradigmatic 3-SAT problem, it is plausible that a similar exponential scaling advantage holds for other problems which are naturally described as a PUBO, such as those exemplified in section III. In the future, it will be interesting to quantitatively study potential speedups also for these and other problems. Another interesting question is whether the presence of multi-qubit operators modifies the behavior of quantum resources during the annealing sweep [63, 64, 65, 66]. Finally, the favorable scaling of the adiabaticity time for certain families of 3-SAT problems, as suggested by our analysis of the minimum energy gap in section IV.2.3, encourages more systematic studies comparing the time complexity of quantum annealing to state-of-the-art classical algorithms.
@@ Acknowledgements.
We acknowledge useful discussions with Ivan Boldin, Michael Möller, Junichi Okamoto, Sebastian Rubbert, Marcel Seelbach Benkner, Theerapot Sriarunothai, and Christof Wunderlich. The work reported in this publication is based on a project that was funded by the German Federal Ministry for Education and Research under the funding reference number 13N16437. The authors are solely responsible for the content of this publication. This work has benefited from Q@TN, the joint lab between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. We acknowledge support by Provincia Autonoma di Trento.

### v3

@@ IV.1.3 PUBO-to-QUBO reduction
To assess the annealing performance gains of PUBO over QUBO formulations when solving 3-SAT instances, we compare the minimum energy gap of the PUBO formulation with that of the equivalent QUBO reduction. Finding an optimal QUBO reduction for a given PUBO problem is in general $\mathsf{NP}$-hard. For the small system sizes considered here, we employ the following brute-force algorithm: replace the pair of variables that appears most frequently in the third-order terms of the PUBO cost function by a slack variable (see section III.1), then continue with the second-most-frequent pair of variables and so on, until all third-order terms have been replaced by quadratic ones with associated slack variables and constraints [44]. For each newly introduced ancilla variable $y=x_{i}x_{j}$, a penalty term $\propto(3y+x_{i}x_{j}-2x_{i}y-2x_{j}y)$ needs to be added to the cost Hamiltonian (see section III.1) and scaled by a suitable factor, thus ensuring that violating the penalty term in favor of the three-body interactions is never energetically favorable. This factor depends on the problem size and structure: in our numerics, we make the simple choice that for every three-body interaction term $a_{ijk}x_{i}x_{j}x_{k}$ that is to be reduced by the ancilla variable $x_{\mathrm{ancilla}}=x_{i}x_{j}$, we add $|a_{ijk}|+1$ to the strength of the corresponding penalty term. In section IV.2.5, we discuss how varying the strength of the penalty terms relative to the original cost Hamiltonian affects the overall annealing performance.
@@ IV.2.4 Dependence of the minimum energy gap on the QUBO penalty strength
To summarize, our example demonstrates that there is in general potential for optimization of the minimum energy gap in the QUBO reduction of a native PUBO problem by tuning the strength of the penalty terms. However, finding the optimal penalty strength generally requires solving the problem itself, which is prohibitively expensive for large exponentially hard problems. For our scaling analysis of the minimum gap, we have thus chosen the penalty strength in a problem-agnostic way as described in section IV.2.3. This heuristic choice is simple, incurs only moderate performance overhead, and reliably produces feasible solutions. Further optimizations may be achieved through more sophisticated schemes to approximate the optimal penalty strength [32], but such improvements would have only minor effects on the results of our PUBO versus QUBO analysis, as in our scenario the simple approach is already close to optimal.
@@ V Implementing PUBO via multi-qubit interactions in quantum hardware
Figure 6: (a) The three-qubit gate $R_{ZZZ}(\theta)=\exp(-i4\theta\hat{s}_{1}^{z}\hat{s}_{2}^{z}\hat{s}_{3}^{z})$ can be decomposed into four CNOT gates plus a single-qubit rotation. (b) Each CNOT gate can be further decomposed into one $R_{ZZ}$ gate of fixed angle and multiple single-qubit rotations. The quantum gates in the decompositions are defined as $R_{\alpha}(\theta)=\exp(-i\theta\hat{s}^{\alpha})$, with $\alpha\in\{X,Y,Z\}$, $R_{ZZ}(\theta)=\exp(-2i\theta\hat{s}^{z}\hat{s}^{z})$, and $H$ denoting the Hadamard gate.
@@ V.1 Native multi-qubit interactions and synthesis of three-body gates in quantum circuits
Several proposals exist to directly engineer native three-body spin interactions of the type $\hat{s}_{i}^{z}\hat{s}_{j}^{z}\hat{s}_{k}^{z}$ on the hardware level. For example, in trapped-ion systems, these include concurrent off-resonant driving of first and second sideband [47] (also relevant for other three-body interactions [48, 49, 50]) or a pulse sequence using squeezings and displacements of a phonon bus [51]. Another approach is via Floquet engineering of two-body interactions to obtain the three-body terms in a high-frequency expansion [52]. In addition, proposals for effective multi-body interactions on superconducting qubit architectures have been put forward [53]. Most available schemes share the common drawbacks that their implementation is somewhat complex and that the achievable higher-order interaction rates are considerably slower compared to those of two-body interactions. Moreover, the tunability of the relative strengths of three-body interactions between different triples of spins depends on the utilized scheme as well as hardware characteristics. Nonetheless, these techniques may be exploited in next-generation devices to realize analog quantum annealing with higher-order interactions that benefits from the advantages of direct PUBO encodings in terms of spatial and temporal resources.
While native higher-order interactions can also act as physical implementations of multi-qubit gates in digital quantum computers, due to the disadvantages mentioned above, most available platforms implement only single- and two-qubit gates natively. These can realize higher-order interaction gates, required for PUBO formulations, with only polynomial overhead. An example is given in Fig. 6: a three-qubit gate of the form $R_{ZZZ}(\theta)=\exp(-i4\theta\hat{s}_{1}^{z}\hat{s}_{2}^{z}\hat{s}_{3}^{z})$ can be synthesized from four CNOT gates, plus a single-qubit rotation around the $z$-axis [see fig. 6(a)]. Each CNOT gate can be further decomposed into
@@
×$\times R_{y_{1}}\left(-\frac{\pi}{2}\right)R_{x_{1}}\left(-\frac{\pi}{2}\right)R_{x_{2}}\left(-\frac{\pi}{2}\right)H_{1}H_{2}$
×$\times R_{zz}\left(\frac{\pi}{2}\right)H_{1}H_{2}R_{y_{1}}\left(\frac{\pi}{2}\right)\,,$ (17)
where we used a two-qubit controlled rotation gate $R_{zz}$, Hadamard gates $H$ and single-qubit rotations around $x$ and $y$-axis. We show the decomposition in fig. 6(b). The overall circuit for the three-qubit gate then uses four two-body gates of type $R_{zz}$ with fixed angle $\pi/2$ and 29 single-qubit rotations. In this decomposition, two adjacent single-qubit gates could be further combined into only one single-qubit rotation. Similar decompositions are possible with other two-qubit gates like the Mølmer–Sørensen (MS) type gate $R_{xx}$. Note that in digital quantum annealing protocols, the three-body interaction strength in each Trotter step is controlled by the rotation angle $\theta$. Since $\theta$ enters in fig. 6(a) only via a single $R_{Z}(\theta)$ gate, the overall three-qubit gate time is practically the same for different rotation angles.
@@ V.2 PUBO speedup
As suggested by our numerical results in section IV, the minimal gap decreases more slowly with increasing system size for the PUBO formulation of certain 3-SAT problems than for the corresponding QUBO reduction, indicating that there exists a system size above which the PUBO formulation will always be favorable. Specifically, this is the case for 3-SAT instances generated by $\mathsf{toughSAT}$, see fig. 3(b).
@@
Even in scenarios where PUBO and QUBO scale approximately the same, as is the case for the 3-SAT problems generated by $\mathsf{uniquePT1}$ shown in fig. 3(a), a constant PUBO speedup is still possible. In the context of digital quantum annealing (see section II.4), we can use the estimated overhead of the three-body interactions from section V.1 together with the minimal gaps extracted numerically in section IV in order to estimate the potential gain in computing time when using the direct PUBO formulation according to eq. 8. In the example of fig. 6, executing the three-qubit gate involves four two-qubit gates and therefore takes about four times as long as a single two-qubit gate (neglecting single-qubit gate times). In the context of quantum annealing, this corresponds to a ratio of three-body to two-body interaction strength of $J^{(3)}/J^{(2)}\approx 1/4$. Using eq. 8 and the average values for the $\mathsf{uniquePT1}$ generator listed in table 1, we obtain $T_{Q}/T_{P}\approx$12.8$\mathinner{(\tilde{V}_{Q}/\tilde{V}_{P})}\mathrm{e}^{$0.024$\,N}$. In our numerics for small systems, we find a typical ratio of $\tilde{V}_{Q}/\tilde{V}_{P}\approx 0.5$, which only weakly depends on the system size. Our estimate therefore hints at a constant PUBO speedup in this regime; more precisely, an adiabatic sweep can be about six times faster if the PUBO instead of the QUBO formulation for this class of 3-SAT problems is used. If there is on top of that an (even small) exponential enhancement, as suggested by our numerical results, it will then dominate over a possible polynomial scaling of $\tilde{V}_{Q}/\tilde{V}_{P}$ at larger system sizes and further improve the performance of PUBO over QUBO. Note that the precise value of the estimated speedup depends on the choice of the strength of the driving Hamiltonian as well as the strength of the constraint terms in the QUBO reduction. Using a better heuristic for the latter, e.g., the one proposed in Ref. [32], could possibly reduce the calculated PUBO improvement. From our analyses in sections IV.2.4 and IV.2.5, where we found the used parameters to be close to optimal, we expect such a change to be only minor.
@@
For instances generated by $\mathsf{toughSAT}$, we see numerical evidence for a much stronger exponential scaling advantage of the PUBO formulation compared to the corresponding QUBO reduction: under the same assumptions as above, we find $T_{Q}/T_{P}\approx$0.81$\mathinner{(\tilde{V}_{Q}/\tilde{V}_{P})}\mathrm{e}^{$0.49$\,N}$. Already for small systems with $N=11$ classical variables, the adiabaticity time reduces by about two orders of magnitude in the PUBO formulation. That is, solving this family of exponentially hard 3-SAT problems via quantum annealing is expected to be two orders of magnitude faster when using the PUBO formulation over the conventional approach of applying a QUBO reduction, and this estimate increases by another order of magnitude for every $\approx 4.7$ additional classical variables in the problem.
@@ VI Conclusion
In this work, we have presented a variety of combinatorial problems that are naturally phrased in terms of higher-order polynomial unconstrained binary optimization. Using the PUBO formulation directly instead of the corresponding QUBO reduction—currently the standard choice in quantum annealing—can have a number of advantages: (1) PUBO formulations avoid additional ancillary qubits, thus reducing the spatial resources required. (2) The ancillary qubits required by QUBO reductions are typically subject to consistency constraints, which require large energy scales—a central limiting factor in existing quantum annealing hardware. (3) As we have illustrated numerically at the example of the paradigmatic 3-SAT problem, PUBO formulations can have significantly larger minimum energy gaps, allowing for faster annealing sweeps on both analog and digital platforms. This gain can outweigh the overhead incurred by higher-order interactions, as we have exemplified for digital quantum annealing with three-qubit gates synthesized from universal sets of single- and two-qubit gates. In particular, we find an exponential advantage of PUBO over QUBO in the scaling of the minimum energy gap with the system size for certain classes of problems (in our example, already for small $\mathsf{toughSAT}$ instances with $N=11$ variables, this suggests a potential speedup of about two orders of magnitude). Altogether, implementing optimization problems in quantum annealers as native PUBO rather than QUBO can thus lead to savings in terms of both spatial and temporal resources.
@@
Although in this work we have focused on the paradigmatic 3-SAT problem, it is plausible that a similar exponential scaling advantage holds for other problems which are naturally described as a PUBO, such as those exemplified in appendix A. In the future, it will be interesting to quantitatively study potential speedups also for these and other problems. Another interesting question is whether the presence of multi-qubit operators modifies the behavior of quantum resources during the annealing sweep [56, 57, 58, 59]. Finally, the favorable scaling of the adiabaticity time for certain families of 3-SAT problems, as suggested by our analysis of the minimum energy gap in section IV.2.3, encourages more systematic studies comparing the time complexity of quantum annealing to state-of-the-art classical algorithms [17, 18, 19].
@@ Acknowledgements.
We acknowledge useful discussions with Marcel Seelbach Benkner, Ivan Boldin, Michael Möller, Junichi Okamoto, Sebastian Rubbert, Theerapot Sriarunothai and Christof Wunderlich. The work reported in this publication is based on a project that was funded by the German Federal Ministry for Education and Research under the funding reference number 13N16437. The authors are solely responsible for the content of this publication. This work has benefited from Q@TN, the joint lab between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. We acknowledge support by Provincia Autonoma di Trento. This Accepted Manuscript is available for reuse under a CC BY-NC-ND licence after the 12 month embargo period provided that all the terms and conditions of the licence are adhered to.
Appendix A Selected PUBO problems
There exists a plethora of optimization problems that are naturally posed as PUBO. In this appendix, we mention a few relevant examples.
A.0.1 $p$-spin model
The cost Hamiltonian of the $p$-spin model ($p\geq 2$) reads
A.0.2 Hypergraph coloring
A.0.3 NAE-k-SAT
Apart from $k$-SAT, which we have discussed in detail in section III.2, there are many other variations and constraints on the general satisfiability problem. One of those is the NAE-$k$-SAT (Not-All-Equal-k-SAT) problem [70], which can also be naturally formulated as a PUBO. Like $k$-SAT, NAE-$k$-SAT consists of a conjunction of clauses, where each clause contains $k$ Boolean variables and is only true, if at least one variable is true and at least one variable is false. Only for $k\geq 4$ is the problem naturally formulated as a PUBO with higher-order interactions [70]. For example, for $k=4$, a clause $C(x_{i},x_{j},x_{k},x_{l})$ can be encoded into one four-body interaction $\sim\hat{s}^{z}_{i}\hat{s}^{z}_{j}\hat{s}^{z}_{k}\hat{s}^{z}_{l}$ and six two-body interaction penalty terms of the cost Hamiltonian.
A.0.4 Traveling salesperson problem with time windows
Appendix B QUBO formulations of 3-SAT
The most resource-efficient way (in terms of qubit requirements) to formulate a 3-SAT instance as a QUBO problem is the direct reduction of the PUBO formulation to QUBO via the introduction of slack variables (see section III.2.2). Here, we also list two more resource-intensive methods.
B.0.1 3-SAT as QUBO via the maximum independent set problem
A well-known example of a QUBO formulation of 3-SAT uses the equivalent maximal independent set problem (MIS) [72]: Draw a graph with three vertices for each clause (corresponding to the three involved literals). All vertices belonging to the same clause as well as pairs of vertices corresponding to mutually exclusive literals (i.e., $x_{i}$ and $\lnot{x_{i}}$) are connected by edges, the set of which is denoted with $E$. The MIS problem then asks whether it is possible to color $M$ vertices without two colored vertices being connected by an edge. If the answer is yes, the original 3-SAT problem is satisfiable. This MIS problem can be easily translated to a QUBO problem requiring $3M$ qubits (or about $\operatorname*{round}(3\times 4.24\,N)=\operatorname*{round}(12.72\,N)$ for hard problems):
$f_{\mathrm{MIS}}(\bm{x})=\sum\limits_{ij\in E}x_{i}x_{j}\,.$ (20)
Although generically more resources are required than for the reduction with slack variables, a potential advantage of the MIS formulation, especially regarding its use on physical quantum devices, lies in the fact that all interactions have the same energy scale, whereas slack variables for a direct QUBO reduction are introduced with constraints that are imposed through high energy penalties.
B.0.2 3-SAT as QUBO via linear inequality constraints
$\text{max}\sum_{m}z_{m}$ (21)
under the inequality constraints